A little help please!?...

Find the two x-intercepts of the graph of:
g(x)=2x^2-8x+6
*Use the answer format of "x=#, x=#"
Thank you in advance! :)

Answer by stanbon(75887)

Find the two x-intercepts of the graph of:
g(x)=2x^2-8x+6
---
Solve 2x^2-8x+6 = 0
x^2-4x+3 = 0
(x-3)(x-1) = 0
----
x-intercepts are x = 3 and x = 1
===================================
Cheers,
Stan H.